# Dual Conditional Diffusion Models for Sequential Recommendation

### v1

@@ Abstract.
Recent advancements in diffusion models have shown promising results in sequential recommendation (SR). However, current diffusion-based methods still exhibit two key limitations. First, they implicitly model the diffusion process for target item embeddings rather than the discrete target item itself, leading to inconsistency in the recommendation process. Second, existing methods rely on either implicit or explicit conditional diffusion models, limiting their ability to fully capture the context of user behavior and leading to less robust target item embeddings. In this paper, we propose the Dual Conditional Diffusion Models for Sequential Recommendation (DCRec), introducing a discrete-to-continuous sequential recommendation diffusion framework. Our framework introduces a complete Markov chain to model the transition from the reversed target item representation to the discrete item index, bridging the discrete and continuous item spaces for diffusion models and ensuring consistency with the diffusion framework. Building on this framework, we present the Dual Conditional Diffusion Transformer (DCDT) that incorporates the implicit conditional and the explicit conditional for diffusion-based SR. Extensive experiments on public benchmark datasets demonstrate that DCRec  outperforms state-of-the-art methods.
@@
copyright: nonecopyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06
@@ 1. Introduction
Sequential recommender systems (SRSs) are designed to produce recommendations by predicting the next item that will capture a user’s interest (Wang et al., 2019a). Despite their effectiveness, existing models represent items with fixed vectors, which limit their ability to capture the latent aspects of items as well as the diverse and uncertain nature of user preferences (Li et al., 2024b; Yang et al., 2023b). Besides, these methods assume that the items with the most user interactions are the most relevant, which puts less-interacted items at a disadvantage and leads to the exposure problem (Yang et al., 2023b). Diffusion models (Ho et al., 2020; Ho and Salimans, 2022a), as an emergent deep generative model, add random noise to input data in a forward process and then recover the original data through a step-by-step reverse denoising process. Their denoising nature aligns well with the SR task, as modeling sequential item recommendations mirrors this step-wise approach (Xie et al., 2024). Recent studies on diffusion models for SR (Yang et al., 2023b) follow the paradigm of diffusion models, perturbing the embedding of the target item through a forward diffusion process into a Gaussian prior distribution. Then, they restore the Gaussian distribution iteratively through a reverse denoising process, also referred to as the sampling phase, to recover target representations and recommend items that are most similar to it.
Although these methods have achieved significant performance, current diffusion models for SR are still undergoing some shortcomings. Firstly, existing approaches (Yang et al., 2023b) commonly conduct conditional diffusion-based recommendations by conditioning the reverse denoising process on user behavior sequences. These methods focus on optimizing the diffusion process for the target item embeddings or generating an oracle item embedding for recommendation (Lin et al., 2024b). However, they do not explicitly model the diffusion process for the discrete target item itself (Gap 1). As illustrated in Figure 1(a), the traditional methods overlook a critical step: mapping the reversed target item representation into the discrete item index space. These methods often determine the recommended item by calculating the similarity (e.g., inner product) between the reversed target item representation and candidate item embeddings, selecting the item with the highest similarity score. While this method works for item ranking, it does not align with the core principles of diffusion models. Diffusion models are inherently probabilistic and continuous, but this step—based on a direct inner product comparison—introduces a deterministic, discrete decision process that breaks the continuous generative flow of diffusion. This inconsistency creates a gap between the optimization directions of the ranking loss for the recommendation task and the denoising loss for the diffusion model, resulting in suboptimal recommendation performance.
Figure 1. Figures (a) and (b) illustrate the differences between previous methods and our proposed approach in Markov modeling, while Figures (c) and (d) highlight the contrast between implicit and explicit conditional diffusion models for SR.
@@
More importantly, existing diffusion-based SR methods can be broadly classified into two categories: implicit and explicit conditional approaches. As shown in Figure 1(c) and (d), in this paper, Implicit conditional methods refer to treating the user’s historical behavior as an implicit feature and integrating them into the generating target item at each diffusion step (Yang et al., 2023b; Li et al., 2024a, b). In contrast, Explicit conditional methods refer to leverage those historical behaviors as explicit guidance that directly influences the denoising process in diffusion SR (Lin et al., 2024a). The main difference between the two types of conditioning is that the implicit one can be regarded as a feature fusion with fixed and compressed historical features while the explicit one iteratively interacts with the generating target according to the diffusion progress. However, both approaches have inherent limitations. Implicit conditioning may oversimplify complex user behaviors by compressing them into a single vector, potentially losing valuable sequential information. Explicit conditioning, while providing detailed conditioning, can be more sensitive to noise and irrelevant actions within the user’s behavior sequence, which may hinder the recommendation quality (He and McAuley, 2016a). As no existing work combines implicit and explicit conditioning to leverage their respective strengths for enhanced SR (Gap 2). Another problem for existing diffusion models-based SR is their inefficient and slow sampling process, which relies on long Markov chains. Generating a target item requires iteratively refining random noise through numerous denoising steps, making the process computationally intensive and time-consuming (Lin et al., 2024b).
@@
Aiming at bridging the aforementioned two significant gaps in existing SRSs, we propose the Dual Conditional Diffusion Models for SR (DCRec). Firstly, we begin with introducing the discrete-to-continuous sequential diffusion framework for SR. Our approach maintains consistency with the existing diffusion framework by introducing a Markov chain to model the transition from the reversed target item representation to the discrete item index. This probabilistic step fits more naturally within the diffusion process, allowing us to better capture the stochastic nature of item generation and selection. By explicitly modeling this transition, our method not only aligns with the underlying diffusion model but also provides a more coherent and accurate recommendation process. Furthermore, instead of following the existing implicit conditional or explicit conditional diffusion model-based SR methods, we propose novel conditioned diffusion models that integrate both implicit and explicit conditioning strategies. Based on our sequential recommendation diffusion framework, we introduce the Dual Conditional Diffusion Transformer (DCDT). This model concatenates historical information as an implicit condition while leveraging a cross-attention mechanism to incorporate explicit control signals throughout the diffusion process. By embedding dual conditions into both the forward and reverse diffusion stages, our approach ensures more precise and contextually relevant recommendations. In addition, the dual conditional mechanism leads to improved model performance, allowing the inference process to achieve optimal results with only a few sampling steps. This reduction in sampling steps significantly lowers computational overhead, making our model more efficient and better suited for real-time applications. The main contributions of this work are threefold:
  • We propose the Dual Conditional Diffusion Models for the sequential recommendation, introducing a novel diffusion framework for the SR. We explicitly introduce the discrete target items into the continuous diffusion process, defining the unified optimization target that considers both the ranking and the diffusion task.
@@
  • Based on this framework, we propose a unified module, named Dual Conditional Diffusion Transformer (DCDT), which can generate accurate target item embedding based on both implicit and explicit conditions. Furthermore, our framework simplifies the diffusion process by enabling few-step inference instead of the traditional full-step inference.
@@
  • We conduct extensive experiments on various public benchmark datasets and show that DCRec  outperforms the state-of-the-art SR models.
2. Related Work
$\bullet$ Diffusion Models. Diffusion models, derived from non-equilibrium thermodynamics, have achieved notable success in areas like computer vision, sequence modeling, and audio processing (Yang et al., 2024). Two primary paradigms exist unconditional models and conditional models. Unconditional models, such as DDPM (Ho et al., 2020), aim to push the performance boundaries of generative models. Conditional models incorporate guidance from labels to control generated samples. SGGM (Yang et al., 2022) conditions on self-produced labels, while You et al. (You et al., 2023) demonstrate how few labels and pseudo-training enhance both large-scale diffusion models and semi-supervised learners. Dhariwal (Dhariwal and Nichol, 2021) uses classifier guidance to improve sample quality, and Ho and Salimans (Ho and Salimans, 2022b) combine conditional and unconditional scores to balance sample quality and diversity. Multi-modal guidance further enriches the diffusion process, as seen in DiffuSeq (Gong et al., 2023) for NLP tasks and SDEdit (Meng et al., 2022) for image translation, with latent diffusion models (LDM) (Rombach et al., 2022) offering flexible, unified semantics. unCLIP (Ramesh et al., 2022) and ConPreDiff (Yang et al., 2023a) integrate CLIP latents in text-to-image generation.
$\bullet$ Diffusion Model-based Sequential Recommendation. Diffusion models show promise in the sequential recommendation by using user behavior sequences as guidance in the denoising process (Xie et al., 2024; Yang et al., 2023b; Li et al., 2024b). DiffuRec (Li et al., 2024b) corrupts target item embeddings with Gaussian noise and uses a Transformer to iteratively reconstruct them. DreamRec (Yang et al., 2023b) generates the oracle next-item embeddings based on user preferences but struggles with scalability due to its lack of negative sampling. DCDR (Lin et al., 2024a) adopts a discrete diffusion framework for progressive re-ranking. However, current methods face two main issues: they focus on embedding diffusion rather than discrete item generation, leading to inconsistencies, and they rely on implicit or explicit conditioning, limiting the capture of user context and weakening target item embeddings.
@@ 3.1. Problem Formulation
3.2. Diffusion Models for Continuous Domains
Recently, diffusion models have already demonstrated significant performance in continuous domains like computer vision and audio (Ho et al., 2020; Ho and Salimans, 2022a; Kong et al., 2021; Rombach et al., 2022). Thus, before presenting our diffusion SR model, we briefly recap the diffusion models in continuous domains (Ho et al., 2020; Luo, 2022).
@@ 3.2. Diffusion Models for Continuous Domains
A generative model is expected to match estimation distribution $p_{\theta}(e_{0})$ to the ground truth distribution $p(e_{0})$ as closely as possible, where $\theta$ represents learnable parameters and $e_{0}\in\mathbb{R}^{d}$ is a set of samples. For vanilla continuous diffusion models, they model the sample $e_{0}$ as a Markov chain with a sequence of latent variables $\{e_{T},\dots,e_{t},e_{t-1},\dots,e_{0}\}$, where $T$ is the diffusion step and $e_{T}$ is a standard Gaussian, $e_{t}$ is the intermediate latent variable. By this means, the latent dimension is exactly equal to the data dimension of $e_{0}$, and the diffusion model progressively recovers $e_{0}$ from Gaussian noise. Each denoising transition $e_{t}\rightarrow e_{t-1}$ is parameterized by a linear Gaussian transition, calculated by a denoising DNN $f_{\theta}(e_{t},t)$:
(1) $p_{\theta}(e_{t-1}|e_{t})=\mathcal{N}(e_{t-1};\mu_{\theta}(e_{t},t),\Sigma_{\theta}(e_{t},t))$
Thus, the joint distribution of diffusion models can be written as:
(2) $\small p_{\theta}(e_{0:T})=p_{\theta}(e_{T})\prod^{T}_{t=1}p_{\theta}(e_{t-1}|e_{t})$
(3) $where,p_{\theta}(e_{T})=\mathcal{N}(e_{T};\mathbf{0},\mathbf{I})$
@@
To train the denoising model $f_{\theta}(e_{t},t)$, diffusion models first construct a progressive noising process $\{e_{0},\dots,e_{T-1},e_{T}\}$, known as the forward process. Corresponding to this, the denoising process is called the reverse process. The forward process adds Gaussian noise to $e_{0}$ step-by-step until the final latent $e_{T}$ is guaranteed to a standard Gaussian. Each noising transition $e_{t-1}\rightarrow e_{t}$ is defined as a linear Gaussian transition $q(e_{t}|e_{t-1})=\mathcal{N}(e_{t};\sqrt{\alpha_{t}}e_{t-1},(1-\alpha_{t})\mathbf{I})$, where the Gaussian parameters $\alpha_{t}\in(0,1)$ varies over time step $t$ and controls the noise scale added to the embedding in the current step $t$. Due to the Markov property and the Bayes’ theorem, the transition can be rewritten with $e_{0}$ as:
(4) $q(e_{t}|e_{t-1})=q(e_{t}|e_{t-1},e_{0})=\frac{q(e_{t-1}|e_{t},e_{0})q(e_{t}|e_{0})}{q(e_{t-1}|e_{0})}$
To optimize the parameter $\theta$, the diffusion model is trained to maximize the log-likelihood of $p_{\theta}(x)$ according to observed samples $e_{0}$. This objective can be formulated as $\arg\max_{\theta}\log p_{\theta}(e_{0})$. Mathematically, we can derive a proxy objective called the Evidence Lower Bound (ELBO), a lower bound of the likelihood term $\log p_{\theta}(e_{0})$. Formally, the equation of the ELBO is:
(5) $\mathbb{E}_{q(e_{1:T}|e_{0})}\left[\log\frac{p_{\theta}(e_{0:T})}{q(e_{1:T}|e_{0})}\right]$
Rather than directly maximize the likelihood term, diffusion models minimize the minus ELBO as (Luo, 2022; Ho et al., 2020):
(6) $\arg\max\limits_{\theta}\log p_{\theta}(e_{0})$ $\approx\arg\max\limits_{\theta}\mathbb{E}_{q(e_{1:T}|e_{0})}\left[\log\frac{p_{\theta}(e_{0:T})}{q(e_{1:T}|e_{0})}\right]$
(7) $=\arg\min\limits_{\theta}\mathbb{E}_{q(e_{1:T}|e_{0})}\left[\log\frac{q(e_{1:T}|e_{0})}{p_{\theta}(e_{0:T})}\right]$
Combining with Equations 2 and 4, the log term in Equation 7 can be further derived as:
(8) $\log\frac{q(e_{T}|e_{0})}{p_{\theta}(e_{T})}+\sum_{t=2}^{T}\log\frac{q(e_{t-1}|e_{t},e_{0})}{p_{\theta}(e_{t-1}|e_{t}))}-\log p_{\theta}(e_{0}|e_{1})$
Therefore, based on Equations 7 and 8, the ELBO-based optimization object of parameter $\theta$ is to minimize:
@@
$-\log p_{\theta}(e_{0}|e_{1})\Big{]}$
To simplify and combine the second and third terms in Equation 9, recent research (Luo, 2022) derivative a simple surrogate objective to obtain a mean-squared error term:
(10) $\mathcal{L}_{\text{elbo}}(e_{0})=\mathbb{E}_{q(e_{1:T}|e_{0})}\left[\sum_{t=1}^{T}\|f_{\theta}(e_{t},t)-e_{0}\|^{2}\right]$
4. Dual Conditional Diffusion Models for Sequential Recommendation
Figure 2. (a) illustrates a simplified workflow of our DCRec. During the end-to-end diffusion process, DCRec  implicitly concatenates the noised target item embedding $\bm{e_{t}}$ and the corresponding noised history embedding sequence $\bm{H_{t}}$ at each diffusion step. Meanwhile, DCRec  is also guided by clear history embedding $\bm{H_{0}}$ as explicit conditional signals. (b) depicts the forward and reverse diffusion process with conditioning. There are two denoising trajectories with implicit guidance. The dual conditional trajectory with additional explicit guidance is more likely to reach the target item (the orange area).
In this section, we present the comprehensive introduction of our DCRec  as shown in Figure 2, which includes both the Sequential Recommendation Diffusion Framework and the Dual Conditional Diffusion Transformer (DCDT). Firstly, we expend the vanilla continuous diffusion to the discrete sequential recommendation task and provide a theoretical-grounded diffusion framework. Then, we introduce the DCDT, our denoising network $f_{\theta}$ for diffusion SR under the proposed framework. Finally, we focus on the efficient inference for the DCRec.
4.1. Sequential Recommendation Diffusion Framework
Continuous diffusion models have achieved remarkable success in domains like vision and audio (Kong et al., 2021), but their application to SR has been limited due to the inherently discrete nature of the recommended items. Existing diffusion models in SR typically determine the recommended item by computing the similarity (e.g., inner product) between the reversed target item representation and candidate item embeddings, selecting the item with the highest similarity score. While this approach is effective for ranking, it fundamentally conflicts with the principles of diffusion models. Vanilla diffusion models are designed to be probabilistic and continuous. Current diffusion SR approaches (Yang et al., 2023b; Li et al., 2024b; Xie et al., 2024) directly apply the continuous diffusion framework to the SR task neglecting the discrete-to-continuous transition. This breaks the transition Markov chain, the generative flow that diffusion models rely on, creating a disconnect between the ranking objective and the diffusion model’s denoising process. Specifically, this misalignment leads to an inconsistency between the optimization directions of the ranking loss for recommendation and the denoising loss of the diffusion process, resulting in sub-optimal recommendation performance.
To address this issue and motivated by diffusion models in text domains (Li et al., 2022), we explicitly extend continuous diffusion models to discrete item domains as shown in Figure 2(a). As for discrete target item $z$ from the item pool $\mathbb{Z}$, the Markov chain in the forward and reverse processes are extending as $\{z,\dots,e_{T-1},e_{T}\}$ and $\{e_{T},\dots,e_{0},z\}$. Specifically, to map the discrete variables into continuous domains, we define a learnable embedding function $\text{Emb}(z)$. The forward transition process of $z$ is defined as:

### v2

@@ Abstract.
Recent advancements in diffusion models have shown promising results in sequential recommendation (SR). Existing approaches predominantly rely on implicit conditional diffusion models, which compress user behaviors into a single representation during the forward diffusion process. While effective to some extent, this oversimplification often leads to the loss of sequential and contextual information, which is critical for understanding user behavior. Moreover, explicit information, such as user-item interactions or sequential patterns, remains underutilized, despite its potential to directly guide the recommendation process and improve precision. However, combining implicit and explicit information is non-trivial, as it requires dynamically integrating these complementary signals while avoiding noise and irrelevant patterns within user behaviors. To address these challenges, we propose Dual Conditional Diffusion Models for Sequential Recommendation (DCRec), which effectively integrates implicit and explicit information by embedding dual conditions into both the forward and reverse diffusion processes. This allows the model to retain valuable sequential and contextual information while leveraging explicit user-item interactions to guide the recommendation process. Specifically, we introduce the Dual Conditional Diffusion Transformer (DCDT), which employs a cross-attention mechanism to dynamically integrate explicit signals throughout the diffusion stages, ensuring contextual understanding and minimizing the influence of irrelevant patterns. This design enables precise and contextually relevant recommendations. Extensive experiments on public benchmark datasets demonstrate that DCDRec significantly outperforms state-of-the-art methods in both accuracy and computational efficiency.
@@
copyright: nonecopyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Recommender systems
@@ 1. Introduction
Sequential Recommender Systems (SRSs) are designed to produce recommendations by predicting the next item that will capture a user’s interest (Wang et al., 2019a). Despite effectiveness, existing models represent items with fixed vectors, which limit their ability to capture the latent aspects of items as well as the diverse and uncertain nature of user preferences (Li et al., 2024b; Yang et al., 2023b). Besides, these methods assume that the items with the most user interactions are the most relevant, which puts less-interacted items at a disadvantage and leads to the exposure problem (Yang et al., 2023b, 2024a). Diffusion models (Ho et al., 2020; Ho and Salimans, 2022a), add random noise to input data in a forward process and then recover the original data through a step-by-step reverse denoising process. Their denoising nature aligns well with the SR task, as this step-wise approach can mimic the modeling sequential item recommendations (Xie et al., 2024). Recent studies on diffusion models for SR (Yang et al., 2023b; Li et al., 2024b) follow the paradigm of diffusion models, perturbing the embedding of the target item through a forward process into a Gaussian prior distribution. Then, they restore the Gaussian distribution iteratively through a reverse denoising process, also referred to as the sampling phase, to recover target representations and recommend items that are most similar to it. Diffusion-based sequential recommendation differs from standard diffusion in computer vision tasks, as recommender systems aim to fulfill user interests. The generated target should elicit positive user feedback, making it essential for the diffusion model to be more controllable with respect to user preferences.
Figure 1. Figures (a) and (b) are examples of implicit and explicit conditional diffusion models for SR.
@@
In this paper, we broadly classify diffusion-based SR into two categories: implicit and explicit conditional approaches, as shown in Figure 1(a) and (b). Implicit conditional methods refer to treating the user’s historical behavior as an implicit feature and integrating them into the generating target item at each diffusion step (Yang et al., 2023b; Li et al., 2024a, b; Xie et al., 2024). In contrast, Explicit conditional methods refer to leveraging those historical behaviors as explicit guidance that directly influences the denoising process in diffusion SR. The main difference between these two types of conditioning is that the implicit one can be regarded as feature fusion with historical features while the explicit one iteratively interacts with the generating target according to the diffusion progress. Currently, most diffusion-based SR methods are implicit conditional approaches (Yang et al., 2023b; Li et al., 2024a, b; Xie et al., 2024). While promising to some extent, implicit conditioning may oversimplify complex user behaviors by compressing them into a single vector, potentially losing valuable sequential patterns and contextual information. Additionally, explicit conditioning, which offers more precise signals, has shown its potential for directly guide the recommendation process in reranking tasks (Lin et al., 2024a), though it remains unexplored in the sequential recommendation. Moreover, there is no existing work that combines implicit and explicit conditioning to leverage their respective strengths, which is non-trivial for providing more precise recommendations based on user preference.. It requires dynamically integrating these complementary signals to fulfill the diffusion optimization paradigm while avoiding noise and irrelevant patterns within user behaviors. Another problem for existing diffusion models-based SR is their inefficient and slow sampling process, which relies on long Markov chains. Generating a target item requires iteratively refining random noise through numerous denoising steps, making the process computationally intensive and time-consuming (Lin et al., 2024b).
@@
We propose the Dual Conditional Diffusion Model for Sequential Recommendation (DCRec). Specifically, instead of employing a single condition, we propose a novel conditioned diffusion model framework that integrates both implicit and explicit conditioning strategies. Such an approach comprehensively accounts for both explicit and implicit advantages while preserving the evidence lower bound (ELBO) consistency under the diffusion framework. Based on our sequential recommendation diffusion framework, we introduce the Dual Conditional Diffusion Transformer (DCDT). This module concatenates historical information and the expected target item as an implicit condition while leveraging a cross-attention mechanism to incorporate explicit control signals dynamically throughout the diffusion process. By embedding dual conditions into both the forward and reverse diffusion stages, our approach ensures more precise and contextually relevant recommendations. In addition, the dual conditional mechanism leads to improved model performance, allowing the inference process to achieve optimal results with only a few sampling steps. This reduction in sampling steps significantly lowers computational overhead, making our model more efficient and better suited for online applications. The main contributions of this work are threefold:
  • We first propose Dual Conditional Diffusion Models for the sequential recommendation, introducing a novel diffusion framework for SR.
@@
  • Based on this framework, we propose a unified module, Dual Conditional Diffusion Transformer (DCDT), which can generate accurate target item embeddings based on both implicit and explicit conditions. Furthermore, our framework accelerates the diffusion inference process by significantly reducing the sampling steps.
@@
  • We conduct extensive experiments on various public benchmark datasets and show that DCRec  outperforms state-of-the-art SR models. Comprehensive empirical analyses and ablation studies further demonstrate the effectiveness of the proposed framework.
@@ 2. Preliminary
Before presenting our model, we define the scope of the problem in sequential recommendation. Subsequently, we briefly outline the theoretical foundation of vanilla diffusion models for continuous spaces and current diffusion-based sequential recommender models for discrete spaces.
@@ 2.1. Problem Formulation
Figure 2. (a) illustrates a simplified workflow of our DCRec. During the end-to-end diffusion process, DCRec  implicitly concatenates the noisy target item embedding $\bm{e_{t}}$ and the corresponding noisy history embedding sequence $\bm{H_{t}}$ at each diffusion step. Meanwhile, DCRec  is also guided by clear history embedding $\bm{H_{0}}$ as explicit conditional signals. (b) depicts the forward and reverse diffusion process with conditioning. There are two denoising trajectories with implicit guidance. The dual conditional trajectory with additional explicit guidance is more likely to reach the target item (the orange area).
2.2. Diffusion Models
In this section, we first recap the paradigm of vanilla diffusion models for continuous domains like computer vision and audio (Ho et al., 2020; Ho and Salimans, 2022a; Kong et al., 2021; Rombach et al., 2022). Then, we briefly introduce the expansion of diffusion models for discrete sequential recommendation task.
2.2.1. Vanilla Diffusion Models
@@
Diffusion Models is expected to match estimation distribution $p_{\theta}(e_{0})$ to the ground truth distribution $p(e_{0})$ as closely as possible, where $\theta$ represents learnable parameters and $e_{0}\in\mathbb{R}^{d}$ is a set of samples. They recover the sample $e_{0}$ through a progressive denoising process modeled as a Markov chain $\{e_{T},e_{T-1},\dots,e_{0}\}$, where $T$ is the diffusion step and $e_{T}$ is a standard Gaussian, $e_{t}$ is the intermediate latent variable. Each denoising transition $e_{t}\rightarrow e_{t-1}$ is parameterized by a linear Gaussian transition. The arithmetic mean $\mu_{\theta}$ and standard deviation $\Sigma_{\theta}$ are calculated by a denoising DNN $f_{\theta}(e_{t},t)$. Therefore, the transition with trainable parameters $\theta$ can be formulated as follows:
(1) $p_{\theta}(e_{t-1}|e_{t})=\mathcal{N}(e_{t-1};\mu_{\theta}(e_{t},t),\Sigma_{\theta}(e_{t},t)),$
@@
Diffusion models train the denoising model $f_{\theta}(e_{t},t)$ by constructing a progressive noising process $\{e_{0},\dots,e_{T-1},e_{T}\}$, known as the forward process, with the aforementioned denoising process referred to as the reverse process. The forward process adds Gaussian noise to $e_{0}$ step-by-step until the final latent $e_{T}$ is guaranteed to a standard Gaussian. Each noising transition $e_{t-1}\rightarrow e_{t}$ is defined as a linear Gaussian transition:
(2) $q(e_{t}|e_{t-1})=\mathcal{N}(e_{t};\sqrt{\alpha_{t}}e_{t-1},(1-\alpha_{t})\mathbf{I}),$
where the Gaussian parameters $\alpha_{t}\in(0,1)$ varies over time step $t$ and controls the noise scale added to the embedding in the current step $t$. The diffusion model is trained to maximize the log-likelihood of $p_{\theta}(x)$ according to observed samples $e_{0}$, which can be formulated as $\arg\max_{\theta}\log p_{\theta}(e_{0})$. Following (Ho et al., 2020), the canonical objective of a diffusion model is to optimize the Evidence Lower Bound (ELBO) of $\log p_{\theta}(e_{0})$:
@@
$-\log p_{\theta}(e_{0}|e_{1})\Big{]},$
To simplify Equation 3, recent research (Luo, 2022) derivatives a simple surrogate objective to obtain a mean-squared error term:
(4) $\mathcal{L}_{\text{elbo}}(e_{0})=\mathbb{E}_{q(e_{1:T}|e_{0})}\left[\sum_{t=1}^{T}\|f_{\theta}(e_{t},t)-e_{0}\|^{2}\right],$
2.2.2. Diffusion Models for Sequential Recommendation
The inherently discrete item domain in sequential recommendation poses challenges for applying standard diffusion models, which are originally designed for continuous domains. Considering discrete target item $z$ from the item pool $\mathbb{Z}$, recent diffusion-based SR approaches (Yang et al., 2023b; Li et al., 2024b; Xie et al., 2024; Du et al., 2023) define an embedding function $\text{Emb}(z)$ that maps discrete items into a continuous space. The transition process of $z$ to latent continuous vectors $e_{0}$ is defined as: